OpenStudy (saadsword):

Need help with probability: If three coins are flipped, what is the probability that there are more heads than tails?

OpenStudy (brainzonly):

*yawns* can't anyone come up with hard questions? i'm bored...

OpenStudy (brainzonly):

anyway P (3 heads) = (1/2) ³ = 1/8 P (3 tails) = (1/2) ³ = 1/8 P (3 heads or 3 tails) = 1/8 + 1/8 - 1/64 = 15/64 Enjoy!

OpenStudy (saadsword):

the book says the answer is 1/2

OpenStudy (brainzonly):

whatever. Use the book then

OpenStudy (saadsword):

maybe you forgot P(2 heads and 1 tails)

OpenStudy (brainzonly):

hey, you're right :D

OpenStudy (brainzonly):

sorry about that. (After you told me about the book I was looking on this 1 site that said the answer I gave you and then it had 1/8ths, 1/4s... it was crazy!!!!!)

OpenStudy (brainzonly):

have a good day

OpenStudy (saadsword):

Thanks

OpenStudy (saadsword):

so its P(3 heads and 2heads)?

OpenStudy (brainzonly):

yes

OpenStudy (saadsword):

1/8 * 1/4?